# The NTU Partitioned Matching Game for International Kidney Exchange Programs

### v1

@@ 2.3 Existence of a strong core matching
To see that the conditions are necessary, observe that $|V(M)\cap A|=2$ must hold for every $A\notin\mathcal{K}^{*}$. Also, if $(A,B)$ is a pair, then $|V(M)\cap(A\cup B)|\geq 3$ must hold by Lemma 9. Lastly, if $A\notin\mathcal{K}_{0}$, then $|V(M)\cap A|\geq 1$ must hold by Proposition 8.
To see sufficiency, we use the alternating path characterization in Lemma 5 for membership in $\mathcal{SC}$. Suppose $M$ satisfies the above two conditions. Then, the properties of $M$ imply that if a player-edge is in an $M_{0}$-alternating cycle, then both of its vertices are covered by $M$. Let $P$ be an $M_{0}$-alternating path from $A\in\mathcal{K}$ to $B\in\mathcal{K}$, where none of the vertices of $A$ are covered by $M$. Then $A\in\mathcal{K}_{0}$, which contradicts that there is an $M_{0}$-alternating path to some $B\in\mathcal{K}$ from $A$. Finally, suppose that $P$ is an $A-B-C$ alternating path, where $|V(M)\cap B|=1$. If $A\notin\mathcal{L}^{*}$, $B\notin\mathcal{L}^{*}$ or $C\notin\mathcal{L}^{*}$, then one of $A,B,C$ is covered twice by $M$, and we are done. Otherwise, since $A,C\in\mathcal{L}^{*}\subseteq\mathcal{K}$ and $B\in\mathcal{L}^{*}\subseteq\mathcal{L}$, the definition of $\mathcal{L}$ implies that $(A,B)$ is a pair or $(B,C)$ is a pair. In the former case, $|V(M)\cap A|=2$, while the latter case implies $|V(M)\cap C|=2$ by the second condition on $M$. Thus, the conditions of Lemma 5 are satisfied and therefore $M$ is in the strong core. ∎
Finding a matching with the properties in Theorem 13 can be done in polynomial time, since it amounts to finding a common base of a matching matroid and a generalized partition matroid. It remains to show that membership in $\mathcal{L}$ and $\mathcal{L}^{*}$ can be decided in polynomial time, which follows from the next lemma.
@@
For given $A,B,C\in\mathcal{K}$, it can be decided in polynomial time if an $(A,B;C)$-$\delta$-path exists.
@@
We consider the graph $G^{\prime}$ obtained from $E\cup M_{0}$ by removing $A$. Since $A\in\mathcal{K}$, there is no $M_{0}$-alternating path between $a$ and $a^{\prime}$, so we can consider the GE decomposition. We may assume that $B$ has a vertex $b$ in some $Q_{i}$ and $C$ has a vertex $c$ in $Q_{j}$ for some $j\neq i$, since otherwise no $(A,B;C)$-$\delta$-path can exist. Since an $(A,B;C)$-$\delta$-path contains an alternating $a-c$ path and an alternating $a^{\prime}-c$-path, we may also assume $j\in[k]$.
Case 1: $i\in\{k+1,k+2\}$.
We may assume $i=k+1$ by symmetry. First, we find an $M_{0}$-alternating path $P_{1}$ from $a^{\prime}$ to $C$ (if no such path exists, then there is no $(A,B;C)$-$\delta$-path). Note that $P_{1}$ is disjoint from $Q_{i}$, and contains $S_{j}$.
Next, we check whether there is an $M_{0}$-alternating $a-s_{j}$ path that passes through $B$. To do this, we consider the graph $G^{\prime\prime}$ obtained from $G^{\prime}$ by adding $as_{j}$ and removing $B$ and $S_{j}$, and we check if there is an $(M_{0}\cup\{as_{j}\})$-alternating path from $b$ to $b^{\prime}$ in $G^{\prime\prime}$. Since no $M_{0}$-alternating cycle contains $B$, such a path must contain $as_{j}$, and therefore it gives an $M_{0}$-alternating path $P_{2}$ from $a$ to $s_{j}$ that traverses $B$. If $P_{2}$ does not exist, then there is no $(A,B;C)$-$\delta$-path.
Observe that $P_{2}$ must intersect $P_{1}$ because both contain $s_{j}$. Let $P^{\prime}_{2}$ be the subpath of $P_{2}$ up to the first intersection with $P_{1}$. Since $P_{1}$ is disjoint from $Q_{i}$, $P^{\prime}_{2}$ passes through $B$. Therefore, $P_{1}\cup P^{\prime}_{2}\cup\{A\}$ is an $(A,B;C)$-$\delta$-path.
Case 2: $i\in[k]$.
We check whether there are two vertex-disjoint $M_{0}$-alternating paths between $\{a,a^{\prime}\}$ and $\{s_{i},s_{j}\}$. We can do this by defining $M^{\prime}_{0}$ as $M_{0}\cup\{s_{i}s_{j}\}\setminus\{S_{i},S_{j}\}$, and checking whether there is an $M^{\prime}_{0}$-alternating path between $a$ and $a^{\prime}$. If no such path exists, then there is no $(A,B;C)$-$\delta$-path. Otherwise, by removing $s_{i}s_{j}$, we get a path $P_{1}$ from $a$ or $a^{\prime}$ to $s_{i}$, and a path $P_{2}$ from the other vertex of $A$ to $s_{j}$.
Next, we check whether there is an $M_{0}$-alternating $s_{i}-s_{j}$ path that starts with $S_{i}$ and passes through $B$. If $B=S_{i}$, then this simply means checking whether there is an alternating $s_{i}-s_{j}$ path that starts with $S_{i}$. Otherwise, we consider the graph $G^{\prime\prime}$ obtained from $G^{\prime}$ by adding $s^{\prime}_{i}s_{j}$ and removing $B,S_{i},S_{j}$, and checking whether there is an $(M_{0}\cup\{s^{\prime}_{i}s_{j}\})$-alternating path from $b$ to $b^{\prime}$ in $G^{\prime\prime}$. Since no $M_{0}$-alternating cycle contains $B$, such a path must contain $s^{\prime}_{i}s_{j}$, and therefore it gives an $M_{0}$-alternating path $P_{3}$ from $s_{i}$ to $s_{j}$ that starts with $S_{i}$ and traverses $B$. If $P_{3}$ does not exist, then there is no $(A,B;C)$-$\delta$-path.
We assume that $P_{3}$ exists. Observe that the only common vertex of $P_{1}$ and $P_{3}$ is $s_{i}$, otherwise $B$ would be in an $M_{0}$-alternating cycle. Also, $P_{3}$ must intersect $P_{2}$ because both end at $s_{j}$. Let $P_{3}^{\prime}$ be the subpath of $P_{3}$ up to the first intersection with $P_{2}$.
Finally, we find an $M_{0}$-alternating path $P_{4}$ from $s_{j}$ to $C$ that starts with $S_{j}$. If no such path exists, then there is no $(A,B;C)$-$\delta$-path. The path $P_{4}$ cannot leave $Q_{j}$, so it intersects $P_{1}\cup P_{2}\cup P^{\prime}_{3}$ only in $s_{j}$. Therefore, $P_{1}\cup P_{2}\cup P^{\prime}_{3}\cup P_{4}\cup\{A\}$ is an $(A,B;C)$-$\delta$-path. ∎
2.4 Examples
We present two examples in Figure 3 in order to explain the notions in the proof. The solid edges are in $M_{0}$, while the dashed edges are in $E$. In both examples, $\mathcal{K}=M_{0}$ because there is no $M_{0}$-alternating cycle, and $\mathcal{K}_{0}=\emptyset$. Also, $B\notin\mathcal{L}$ in both examples, since there is an $A-B-E$ alternating path, but there is neither an $(A,B;E)$-$\delta$-path nor an $(E,B;A)$-$\delta$-path. It can be checked that the remaining players are in $\mathcal{L}$ in both examples.
Figure 3: Solid edges are in $M_{0}$, dashed egdes are in $E$. The example on the right is obtained from the example on the left by adding $ab^{\prime}$. The strong core is empty on the left, while bold edges on the right form a matching in the strong core.
In the left example, the pairs are $(A,D),(A,E),(C,D),(C,E),(D,E)$, which means that =$\mathcal{L}^{*}=\{A,C,F\}$, and each are singletons in $G^{*}$. It can be checked that there is no matching that covers each of $A,C,F$ at least once and the remaining players twice, so $\mathcal{SC}=\emptyset$.
In the right example, $(A,C)$ is also a pair, so =$\mathcal{L}^{*}=\{A,C,D,E,F\}$, where $F$ is a singleton in $G^{*}$, while $A,C,D,E$ form a clique. By Theorem 13, $M\in\mathcal{SC}$ if and only if it covers both vertices of $B$, and it leaves at most one vertex of $F$ and at most one vertex of $A\cup C\cup D\cup E$ exposed. One such matching $M$ is indicated in bold; the dotted path is the single alternating path in $M\Delta M_{0}$.
3 Constant number of players
In this section, we show that if the number of players is bounded by a constant, then both the existence questions of a strong (resp. weak) core solution and the verification versions are solvable in polynomial time. This is relevant for example in the IKEP application, where players correspond to countries, which obviously implies a fairly small upper bound on the number of players. We shall see that the tractability of this case is a direct consequence of Lemma 1.
Recall that the instance is given by a graph $G=(V,E)$ and a partition $V=V_{1}\cup V_{2}\cup\dots\cup V_{m}$, where $m$ is the number of players. First, consider the verification problems. Lemma 1 immediately implies that if $m$ is constant, then it can be decided in polynomial time if a given matching $M$ is in the weak core. Indeed, the number of possible coalitions is polynomial, and for a given coalition $\{i_{1},i_{2},\dots,i_{k}\}$, we can decide, using Lemma 1, if there is a matching $M^{\prime}$ in the induced subgraph $G[V_{i_{1}}\cup V_{i_{2}}\cup\dots\cup V_{i_{k}}]$ such that $|V(M^{\prime})\cap V_{i_{j}}|\geq|V(M)\cap V_{i_{j}}|+1$ for every $j\in[k]$.
Membership in the strong core can be decided similarly, but for every coalition $\{i_{1},i_{2},\dots,i_{k}\}$ and every $\ell\in[k]$, we check if there is a matching $M^{\prime}$ in the induced subgraph $G[V_{i_{1}}\cup V_{i_{2}}\cup\dots\cup V_{i_{k}}]$ such that $|V(M^{\prime})\cap V_{i_{j}}|\geq|V(M)\cap V_{i_{j}}|$ for every $j\in[k]\setminus{\ell}$ and $|V(M^{\prime})\cap V_{i_{\ell}}|\geq|V(M)\cap V_{i_{\ell}}|+1$.
@@ 3 Constant number of players
If the number of players is bounded by a constant $m$, then it can be decided in polynomial time whether a matching $M$ is in the weak (resp. strong) core.
Deciding non-emptiness of the weak and strong cores is somewhat more difficult, but still polynomial-time solvable. The crucial observation is that the membership of a matching $M$ in these cores depends only on the values $|V(M)\cap V_{i}|$ ($i\in[m]$). Using Lemma 1, we can find in polynomial time all component-wise maximal vectors $x\in\mathbb{Z}_{+}^{m}$ for which a matching $M$ with $|V(M)\cap V_{i}|\geq x_{i}$ for every $i\in[m]$ exists. Indeed, $x_{i}\leq|V_{i}|$ ($i\in[m]$) can be assumed, so we can check all possible vectors since $m$ is constant. Let $\mathcal{X}$ denote the set of such vectors. For each $x\in\mathcal{X}$, we can check in polynomial time, by using Lemma 1 for every coalition, whether any matching $M$ with $|V(M)\cap V_{i}|\geq x_{i}$ ($i\in[m]$) is in the strong or weak core. If the answer is negative for every $x\in\mathcal{X}$, then the core is empty, otherwise it is non-empty.

### v2

@@ 2.2.1 Deciding the existence of a strong core matching
Suppose that the conditions of the lemma hold, but $u\in A_{L}$ and $P_{ub}\subseteq P^{L}_{A_{L}}\cup P^{L}$ cannot be chosen such that the first case holds, i.e., $P^{L}_{u}\cup P$ is an $M_{0}$-alternating path. We show that in this case, there must be choices of $u\in A_{L}$ and $P_{ub}\subseteq P^{L}_{A_{L}}\cup P^{L}$ such that the second case holds.
Choose $u\in A_{L}$ and $P_{ub}\subseteq P^{L}_{A_{L}}\cup P^{L}$ so that $P=P_{ub}\cup P^{L}_{bc^{\prime}}$ is an $M_{0}$-alternating path and $P^{L}_{u}$ first intersects $P$ (after $u$) as close to $c^{\prime}$ on $P$ as possible. Such a choice must exist because of the existence of $P^{L}$ and because of our assumption that $P$ cannot be chosen to be disjoint from $P_{u}^{L}$.
Let $v$ be this first intersection of $P_{u}^{L}\setminus\{u\}$ with $P$, and let $P^{L}_{uv}$ be the part of $P^{L}_{u}$ from $u$ to $v$. If $v$ is in the part $P^{L}_{bc^{\prime}}$ of $P$ and $v$ is the vertex of its player closer to $b$ on $P$, then $P_{uv}^{L}\cup P$ is an $(A_{L},B_{L};C_{L})$-$\delta$-path $H$ with $P\subseteq H$, and we are done. We show that no other case is possible.
Suppose that $v$ is in the part $P^{L}_{bc^{\prime}}$ of $P$, but $v$ is the vertex of a player that is closer to $c^{\prime}$ on $P$. Then $P_{uv}^{L}\cup P_{uv}$ gives an $M_{0}$-alternating cycle through $A_{L}\in\mathcal{K}$, where $P_{uv}$ is the part of $v$ between $u$ and $v$, contradiction.
Suppose finally that $v$ is in the part of $P$ between $u$ and $b$. Let $P_{vb}$ be the part of $P$ from $v$ to $b$. Since $A_{L}\in\mathcal{K}$, we have that $u^{\prime}:=\{a,a^{\prime}\}\setminus\{u\}$ and $P_{u^{\prime}b}:=\{u^{\prime}u\}\cup P^{L}_{uv}\cup P_{vb}\subseteq P^{L}_{A_{L}}\cup P^{L}$ are also valid choices in the lemma, since if $P_{u^{\prime}b}$ was not an $M_{0}$-alternating path (i.e. if the two edges adjacent to $v$ in $P_{u^{\prime}b}$ were both from $E$), then we would get an $M_{0}$-alternating cycle $P^{L}_{uv}\cup P_{uv}$ through $A_{L}$, a contradiction. Furthermore, as $P^{L}_{uv}\subseteq P^{L}_{A_{L}}$ and $P^{L}_{u^{\prime}}\subseteq P^{L}_{A_{L}}$ are disjoint, we must have that either $P^{L}_{u^{\prime}}$ does not intersect $P^{\prime}=P_{u^{\prime}b}\cup P^{L}_{bc^{\prime}}$, which contradicts our assumption that it is impossible to choose such a path $P_{u^{\prime}b}\subseteq P^{L}_{A_{L}}\cup P^{L}$, or $P^{L}_{u^{\prime}}$ first intersects $P^{\prime}$ in the part $P_{vb}\cup P^{L}_{bc^{\prime}}\setminus\{v\}$. Since this part is the same in $P$ and $P^{\prime}$, we get that this choice of $u^{\prime}$ and $P_{u^{\prime}b}$ leads to a case where the first intersection of $P^{L}_{u^{\prime}}$ with $P^{\prime}$ is closer to $c^{\prime}$ on $P^{\prime}$ than $v$ was to $c^{\prime}$ on $P$, in contradiction to the choices of $u$ and $P_{ub}$. ∎
We remark that if $B_{L}=C_{L}$ in Lemma 9, then the path $P$ given by the lemma always satisfies the first property, because there is no $(A_{L},B_{L};B_{L})$-$\delta$-path.
@@
Let $A,B\in\mathcal{K}$, and let $P$ be an $M_{0}$-alternating ABA-B path. Then for any $M\in\mathcal{SC}$ we have $|V(M)\cap A|\geq 1$ and $|V(M)\cap B|\geq 1$.
@@
Suppose that $M\in\mathcal{SC}$ is a matching where one of them, say $B$ has 0 vertices covered. By Lemma 6 this means that $|V(M)\cap A|=2$. Let $x_{A}$ and $y_{A}$ denote the two endpoints of the path $P_{A}$ containing $e_{A}$ in $M\Delta M_{0}$. The existence of $P$ implies that there is an $M_{0}$-alternating path from $B$ to $P_{A}$, hence also from $B$ to $x_{A}$ or $y_{A}$. Such a path gives a blocking coalition, as $B$ improves and no player gets worse. ∎
Lemma 11.
Consider an $(A,B;C)$-$\delta$-path $H$ with $C\in\mathcal{K}$, and an $M\in\mathcal{SC}$ such that $|V(M)\cap A|=1$ and $|V(M)\cap C|=2$. Then there is an $(A,B;C^{\prime})$-$\delta$-path $H^{\prime}$ with the same cycle as $H$, such that $|V(M)\cap C^{\prime}|\leq 1$ and $C^{\prime}\in\mathcal{K}^{*}$.
Proof.
Let $P_{C}=P_{c}\cup\{cc^{\prime}\}\cup P_{c^{\prime}}$ be the $M\triangle M_{0}$ alternating path through $e_{C}$. Let $P_{dc}=\mathrm{pa}(H)$ be the path of the $\delta$-path $H$ that ends in $c\in C$ and starts in $d\in D$ (where $d$ is the common vertex of the path and the cycle of $H$).
We use Lemma 9 as follows. We set $A_{L}:=C$ and $B_{L}=C_{L}:=D$. The path $P^{L}$ is set to be $P_{dc}$ and $P^{L}_{A_{L}}$ is set to be the maximal subpath of $P_{C}$ that contains $e_{C}$, but does not intersect $\mathrm{cyc}(H)$ (or $\mathrm{cyc}(H)\setminus\{d\}$ if $C=D$). As $e_{C}\notin\mathrm{cyc}(H)$, there is such a subpath. For $u\in\{c,c^{\prime}\}$, let $P^{L}_{u}=P_{u}\cap P^{L}_{A_{L}}$.
Lemma 9 gives us $u\in C$ and $P_{ud^{\prime}}$ such that the path $P_{u}^{L}\setminus\{u\}$ does not intersect $P_{ud^{\prime}}\cup\{dd^{\prime}\}$ nor $\mathrm{cyc}(H)$, by the choice of $P_{A_{L}}^{L}$ (if $C=D$ then $u\neq d$ and $u^{\prime}=d$, so $P_{u}^{L}$ is disjoint from $\mathrm{cyc}(H)$). We claim that $P_{u}$ does not intersect $P_{ud^{\prime}}\cup\{dd^{\prime}\}\cup\mathrm{cyc}(H)$, which can be seen as follows. Suppose for the contrary that there is a nomempty intersection. As, $P^{L}_{u}$ is a subpath of $P_{u}$ (and one that also starts in $u$), which is disjoint from $\mathrm{cyc}(H)\setminus\{d\}$, and $d\in\mathrm{cyc}(H)$, we get that if $P_{u}\setminus\{u\}$ intersects $P_{ud^{\prime}}\cup\{dd^{\prime}\}\cup\mathrm{cyc}(H)$, then by the choice of $P^{L}_{A_{L}}$ the first intersection (from $u$) must be on $\mathrm{cyc}(H)$. Let this intersection point be $v$ and $P_{uv}$ be the part of $P_{u}$ from $u$ to $v$. Then, let $Q_{vd}$ be the subpath of $\mathrm{cyc}(H)$ between $v$ and $d$ that starts with $vv^{\prime}\in M_{0}$, which must exists if $v\neq d$; if $v=d$ we let $Q_{vd}=\emptyset$. Then, we get an $M_{0}$-alternating cycle $P_{uv}\cup Q_{vd}\cup\{dd^{\prime}\}\cup P_{ud^{\prime}}$ ($P_{ud^{\prime}}$ and $Q_{vd}\subseteq\mathrm{cyc}(H)$ are disjoint as $P_{ud^{\prime}}\subseteq(P_{dc}\cup P^{L}_{A_{L}})\setminus\{dd^{\prime}\}$ by Lemma 9) through $C\in\mathcal{K}$, a contradiction.
Hence, we obtain an $(A,B;C^{\prime})$-$\delta$-path by taking the union $\mathrm{cyc}(H)\cup\{dd^{\prime}\}\cup P_{ud^{\prime}}\cup P_{u}$, where $C^{\prime}$ is the last player on $P_{u}$, and it is in $\mathcal{K^{*}}$ because $|V(M)\cap C^{\prime}|\leq 1$.
@@
If there is an $(A,B;C)$-$\delta$-path and an $M\in\mathcal{SC}$ such that $|V(M)\cap A|=1$, then $|V(M)\cap B|=2$.
Proof.
Otherwise, the $(A,B;C^{\prime})$-$\delta$-path given by Lemma 11 would contain a blocking path for $M$ by Lemma 6. ∎
Lemma 13.
Let $Q$ be an $M_{0}$-alternating ABCA-B-C path for some $A\in\mathcal{K}$, $B\in\mathcal{K}^{*}$, $C\in\mathcal{K}$. Then there exists an $(A,B;C)$-$\delta$-path or a $(C,B;A)$-$\delta$-path.
Proof.
Let $M\in\mathcal{SC}$ such that $|V(M)\cap B|=1$ and suppose for the contrary that there exists no $(A,B;C)$-$\delta$-path or $(C,B;A)$-$\delta$-path. Let $P_{A}$ be the $M\triangle M_{0}$ path that contains $e_{A}$ and $P_{C}$ be the $M\triangle M_{0}$ path that contains $e_{C}$. Let the order of vertices on the ABCA-B-C $M_{0}$-alternating path $Q$ be $a,a^{\prime},b,b^{\prime},c,c^{\prime}$.
We use Lemma 9 by setting $A_{L},B_{L},C_{L}$ in the lemma as $A,B,C$ with $M_{0}$-alternating paths $P^{L}:=Q$ and $P^{L}_{A_{L}}:=P_{A}$, which gives us $u\in\{a,a^{\prime}\}$ and $P_{ub}\subseteq P_{A}\cup Q$. Let $Q^{\mathrm{new}}=P_{ub}\cup Q_{bc^{\prime}}$ denote the ABCA-B-C $M_{0}$-alternating path we get by applying the lemma and possibly replacing the part of $Q$ between $a$ and $b$ because of it. As there is no $(A,B;C)$-$\delta$-path, we get that $Q^{\mathrm{new}}$ is disjoint from $P_{u}\setminus\{u\}\subseteq P_{A}$. Note that the part of $Q^{\mathrm{new}}$ and $Q$ between $b$ and $c^{\prime}$ is the same.
Then, we use Lemma 9 again, by setting $A_{L},B_{L},C_{L}$ as $C,B,A$ with $P^{L}:=Q^{\mathrm{new}}$ and $P^{L}_{A_{L}}:=P_{C}$. The conditions of Lemma 9 are satisfied since $C\in\mathcal{K}$. As there is no $(C,B;A)$-$\delta$-path, Lemma 9 gives $w\in C$, and $M_{0}$-alternating paths $P_{b^{\prime}w}\subseteq P_{C}\cup Q^{\mathrm{new}}$ such that $P_{w}\setminus\{w\}\subseteq P_{C}$ is disjoint from $P^{\prime}=Q^{\mathrm{new}}_{ub^{\prime}}\cup P_{b^{\prime}w}$.
We claim that $P_{u}\setminus\{u\}\subseteq P_{A}$ does not intersect $(P^{\prime}\cup P_{w})\subseteq Q^{\mathrm{new}}\cup P_{C}$. By the above, we know that $P_{u}\setminus\{u\}$ is disjoint from $Q^{\mathrm{new}}$. Hence, if $P_{u}\setminus\{u\}$ intersects $P^{\prime}$, then it must be the case that $P_{A}=P_{C}$. If $u\in P_{w}$, then $P_{w}\setminus\{w\}$ is not disjoint from $P^{\prime}$, contradiction. If $P_{w}\subseteq P_{u}$, then $P_{u}\setminus\{u\}$ is not disjoint from $Q^{\mathrm{new}}$, also a contradiction. Hence, the only possibility left is $P_{u}\cap P_{w}=\emptyset$. Then, since the parts of $P^{\prime}$ and $Q^{\mathrm{new}}$ between $u$ and $b^{\prime}$ are the same, $P_{u}\setminus\{u\}$ must first intersect $P^{\prime}\cup P_{w}$ in the part $P^{\prime}_{b^{\prime}w}$ of $P^{\prime}$. If the intersection point $v$ is the vertex of its player that is closer to $w$ on $P^{\prime}$ then $P^{\prime}_{uv}\cup P^{u}_{uv}$ ($P_{uv}^{u}$ is the part of $P_{u}$ between $u$ and $v$) is an $M_{0}$-alternating cycle through $e_{A}$ and otherwise $P^{u}_{uv}\cup P^{\prime}$ is an $(A,B;C)$-$\delta$-path, both are contradictions.
Hence, we obtain that we can create an $M_{0}$ alternating path $P_{u}\cup P^{\prime}\cup P_{w}$, where the endpoints both have only one vertex covered by $M$. As $B$ has both vertices covered in $P_{u}\cup P^{\prime}\cup P_{w}$, this leads to a weakly blocking coalition to $M$, contradiction. ∎